### v1

# Benchmarking Multimodal LLMs on Recognition and Understanding over Chemical Tables

Abstract
Chemical tables encode complex experimental knowledge through symbolic expressions, structured variables, and embedded molecular graphics. Existing benchmarks largely overlook this multimodal and domain-specific complexity, limiting the ability of multimodal large language models to support scientific understanding in chemistry. In this work, we introduce ChemTable, a large-scale benchmark of real-world chemical tables curated from the experimental sections of literature. ChemTable includes expert-annotated cell polygons, logical layouts, and domain-specific labels—including reagents, catalysts, yields, and graphical components—and supports two core tasks: (1) Table Recognition, covering structure parsing and content extraction; and (2) Table Understanding, encompassing both descriptive and reasoning-oriented question answering grounded in table structure and domain semantics. We evaluated a range of representative multimodal models, including both open-source and closed-source models, on ChemTable and reported a series of findings with practical and conceptual insights. Although models show reasonable performance on basic layout parsing, they exhibit substantial limitations on both descriptive and inferential QA tasks compared to human performance, and we observe significant performance gaps between open-source and closed-source models across multiple dimensions. These results underscore the challenges of chemistry-aware table understanding and position ChemTable as a rigorous and realistic benchmark for advancing scientific reasoning  [^1] https://github.com/lqzxt/ChemTable.
1 Introduction
Recent advances in multimodal large language models (MLLMs) have created new opportunities for mining expert knowledge from scientific literature and are increasingly viewed as catalysts for AI-driven scientific discovery [1]. A growing wave of scientific accelerators—such as MLLMs-based OCR, ChatPaper [2], and ChatPDF [3]—demonstrates the potential of MLLMs to streamline literature parsing, summarization, and interactive reading. From a modeling perspective, MLLMs exhibit strong capabilities in semantic understanding and reasoning, making them well-suited for processing the rich multimodal content of scientific documents [4, 5]. Benchmarks like ChartQA [6] and ChartX [7] have begun exploring visual reasoning over scientific figures. Yet, scientific literature remains one of the most semantically dense and domain-specialized corpora, and understanding it serves as both a valuable application and a rigorous testbed for evaluating the limits of MLLMs [8].
While recent benchmarks have primarily focused on figures and charts, tables remain a largely underexplored yet equally critical modality in scientific literature. In chemistry, tables are concise and information-rich representations of experimental setups, reaction conditions, and empirical results. These tables often combine symbolic expressions, structured variables, and graphical elements—posing significant challenges for existing MLLMs. Despite their importance, there is a lack of realistic, domain-specific datasets designed to evaluate the capabilities of MLLMs in scientific table understanding. This gap motivates our work: to build a comprehensive benchmark that captures the structural complexity, semantic richness, and reasoning demands of real-world chemical tables.
Figure 1: Multimodal Complexity of Chemical Tables and Challenges for MLLMs.
Among scientific tables, those found in chemistry pose especially complex challenges that go far beyond standard layout parsing [9]. A typical chemical table encodes rich experimental workflows through dense symbolic notation (e.g., reagents, ligands), domain-specific abbreviations (e.g., “BINAP”, “TFA”), and embedded visual elements such as molecular structures or reaction schemes [10, 11]. As illustrated in Figure 1, a single row often represents a multi-variable configuration—catalyst, ligand, solvent, additive, with quantitative outcomes like yield, selectivity. These tables also rely on implicit conventions (e.g., ratio formats like “$>$19/1”) and footnoted exceptions, making their semantics both subtle and compact. Interpreting such tables requires aligning symbolic, numeric, and visual information in a domain-aware manner, which presents significant challenges for general-purpose MLLMs not trained on scientific representations [8, 9].
To systematically address these challenges, we introduce ChemTable, a high-quality benchmark designed for recognition and understanding over real-world chemical tables. The dataset comprises over 1,300 tables curated from peer-reviewed chemistry literature, spanning diverse reaction types, experimental conditions, and reporting formats. ChemTable supports two core tasks: table Recognition[12]—including structure reconstruction and content extraction—and table Understanding [13, 14], formulated through more than 9,000 QA instances across two categories: (1) descriptive questions, which evaluate a model’s ability to extract key facts; and (2) reasoning questions, which require comparison, attribution, and domain-grounded inference.
To enable scalable and consistent evaluation, all answers follow a short-form format suitable for MLLMs-based automatic grading. We benchmark seven widely-adopted MLLMs, including representative models from both open- and closed-source sources, and observe significant performance gaps across all tasks. Compared to human and expert performance, current MLLMs lag behind—particularly in interpreting symbolic and embedded graphical elements. Our analysis further reveals key insights, including the symbolic understanding gap, the limited transferability of MLLMs to scientific domains, and substantial variation across model families. We release ChemTable, along with evaluation tools, to support future research in multimodal scientific understanding.
2 Related Work
2.1 Existing Benchmarks on Table Recognition.
Existing table recognition benchmarks mainly focus on two tasks: Table Structure Recognition (TSR) and Table Recognition (TR). The widely used ICDAR-2013 [15] dataset includes 257 manually annotated tables for TSR. Despite its small size, it remains a key benchmark due to its annotation quality. Since 2019, large-scale TR datasets [16, 17, 18, 19] have emerged. However, their annotations are programmatically generated and provide only table structures without cell content, limiting their utility in deeper understanding tasks. To address this, FinTabNet [20] and PubTabNet [21] incorporate logical cell locations, enhancing TSR modeling. SciTSR [22] focuses on scientific documents, offering detailed cell content annotations, thus better supporting academic table recognition.
2.2 Existing Benchmarks on Table Understanding.
Several datasets have been proposed for table understanding tasks, including WikiTQ [23] and WikiSQL [24]. These datasets consist of numerous tables extracted from Wikipedia and serve as critical early benchmarks in table understanding research. In recent years, advancements in natural language processing have extended beyond traditional homogeneous tables to incorporate additional modalities. For instance, HybridQA [25] and MMTab [26] introduce multi-modal or semi-structured sources for more complex reasoning. Moreover, domain-specific benchmarks such as FinQA [27] (financial) and SciTab [28] (scientific) address unique challenges by integrating structured tables with related textual content to support complex reasoning tasks.
Table 1: Comparison of Table Understanding Datasets. We use the following shorthand: Dom. Spe. = Domain Specific, Pict. Moda. = Pictorial Modality, Text. Moda. = Textual Modality, Human Writ. = Human Written, LLM Gene. = LLM Generated.
Table Modal Question Source
Name Dom. Pict. Text Human LLM
Spe. Moda. Moda. Writ. Gene.
WikiTQ [23]
WikiSQL [24]
HybridQA [25]
FinQA [27]
SciTab [28]
MMTab [26]
ChemTable
The ChemTable dataset is different from existing benchmarks in several key ways. It is a comprehensive, open-ended dataset focused on table recognition and understanding in chemistry, created to support scientific question-answering systems. Unlike other datasets, its questions combine visual elements, table data, and chemistry knowledge, making it necessary for models to recognize table content and structure, then reason across that information. This setup reflects real tasks that researchers face when drawing conclusions from data. ChemTable also helps evaluate how well multimodal language models can reason, showing their strengths and weaknesses.
3 ChemTable: A Chemical Table Recognition and Understanding Benchmark
We introduce ChemTable, a benchmark that systematically curates chemical research tables, evaluates table recognition capabilities, and advances table-based question answering in the chemical domain.
3.1 Dataset Construction and Annotation
In this section, we outline our methodology for systematically collecting and annotating chemical research tables to construct domain-specific datasets. We first describe the data sources and table types, followed by the introduction of a structured data annotation protocol that seamlessly integrates structural features (e.g., table structural information, text formatting) with chemical elements.
3.1.1 Table Collection
Journal Selection. The journal selection criteria prioritised top-tier chemistry journals (e.g., ACS Catalysis, JACS, Chem, Angewandte Chemie Int. Ed., and Science) to ensure data credibility and disciplinary relevance. Publications from the past five years (2015–2025) were prioritised to balance cutting-edge insights with foundational studies, while DOI links were systematically documented to ensure traceability and facilitate future validation or expansion of the dataset.
Table Categorization Strategy. The dataset comprises six primary table types: (1) condition optimisation tables, (2) substrate screening tables, (3) chemical structure information tables, (4) reaction feature data tables, (5) property/result comparison tables, and (6) data statistics tables. Condition optimisation and substrate screening tables comprise over 50% of the dataset.
Figure 2: Overview of Metadata, Table Recognition, and Table Understanding in ChemTable.
3.1.2 Table Annotation
For table annotation, we formed a dedicated team to identify table titles, annotations, main content, and additional captions [29]. Each element was annotated with pixel coordinates and OCR-verified text. We also encoded the logical structure of tables, such as row-column relationships. To preserve visual meaning, we carefully recorded stylistic features like boldface, italics, and color. For table-based question answering annotation, descriptive questions are directly derived from the detailed annotations of table elements created in the previous step. Simple reasoning questions are generated using GPT-4.1 [30] and then filtered based on difficulty. For more complex reasoning questions and visually descriptive tasks, we manually annotated 2,122 questions with the help of experienced graduate students. These questions focused on specific topics, such as reaction conditions and yields, while allowing a variety of question styles. Annotators were also encouraged to include unanswerable questions caused by missing data, vague references, or incomplete formatting. To ensure accuracy, all questions were verified by human review and model checks. Details can be found in the Appendix.
3.2 Table Recognition
Table recognition is a fundamental step in document understanding, which aims to turn table images into structured data. This task is more challenging in chemistry because of complex layouts and specialized symbols [31].
3.2.1 Task Definition
We employ the generation paradigm of MLLMs to address the table recognition (TR) task, which is formulated as a format mapping problem from images to sequences. Formally, given a dataset $\mathcal{D}=\{(I^{i},H^{i})\}_{i=1}^{n}$ with $n$ samples, we predict the corresponding structured form $H^{i}$ for each table image $I^{i}$. Specifically, we provide the image table $I^{i}$ along with a prompt $P$ as input to the MLLMs, which generates the structured data form $\hat{H}^{i}=\operatorname{MLLM}(P,I^{i})$.
3.2.2 Evaluation Protocols
To further assess the capabilities of table recognition models and understand the challenges of chemical-domain data, we propose a set of tasks focused on domain-specific reasoning in chemical table understanding. Specifically, we introduce the following three tasks:
Value Retrieval: This task evaluates a model’s ability to locate and extract cell-level content accurately. Given a table and a pair of coordinates, the model must return the exact value in that cell. This task directly measures the model’s precision in structured data parsing and positional alignment.
Position Retrieval: This task requires the model to infer the position of a specific value. Given a table and a target value, the model must identify the correct row and column. This tests the model’s understanding of value localisation within structured layouts.
Molecular Recognition: Chemical tables often include molecular structures represented as images, either embedded within cells or positioned externally. This task aims to evaluate a model’s ability to recognise and interpret such molecular graphics. The objective is to extract the corresponding SMILES (Simplified Molecular Input Line Entry System) string from a molecular diagram. This task presents unique challenges, such as fine-grained visual understanding and domain-specific symbol interpretation, which are not typically encountered in general-domain table recognition [32, 33].
Table 2: ChemTable Dataset Statistics. Unique Tokens and Question & Answer Lengths are Calculated Based on the Qwen-7B Tokenizer.
Statistics Value
Images
Total Images $1,382$
Years $2015-2025$
Average size (px) $3687\times 4086$
Descriptive Questions
# questions $7,344$
# unique questions $1,512$
Question
- # unique tokens $1,568$
- maximum length $25$
- average length $11.10$
Answer
- # unique tokens $12,032$
- maximum length $148$
- average length $8.99$
Reasoning Questions
# questions $2,542$
# unique questions $1,735$
Question
- # unique tokens $2,086$
- maximum length $37$
- average length $12.66$
Answer
- # unique tokens $2,610$
- maximum length $78$
- average length $6.21$
3.3 Table Understanding
We divide table comprehension evaluation into two types: descriptive questions and reasoning questions. Descriptive questions test the model’s ability to extract and summarize basic table information, while reasoning questions assess its ability to perform deeper analysis and inference. To improve the quality of the evaluation, we applied a data filtering process to increase both the diversity and difficulty of the questions in our dataset.
3.3.1 Task Definition
We employ the generation paradigm of MLLMs to address the table question answering task. Formally, given samples, each sample consists of a table $T^{i}$, a natural language question $Q^{i}$, and the corresponding answer $A^{i}$. To answer the question, we provide the table $T^{i}$ and the question $Q^{i}$ as input to the MLLMs in the form of a prompt $P$, yielding the predicted answer. In the visual table QA setting, $T^{i}$ is an image of a table, while in the text table QA setting, $T^{i}$ is a structured textual representation.
3.3.2 Descriptive Questions.
Descriptive questions aim to provide a general overview of the basic information presented in the table shown in the image. These questions include: 1) describing the main body of the table, such as its dimensions; 2) describing basic metadata of the table, including titles and notes; 3) describing domain-specific elements, such as reaction conditions in chemical reaction formulas or SMILES notation in molecular graphs; and 4) identifying certain visual features in the table, such as rows highlighted with special colours. Chemical tables often contain images, such as molecular structures, instead of plain text, which makes them harder to read and analyse. For example, an image in a table row can affect spacing and make it harder to recognize content. Sometimes, only certain parts of a molecule, like a red-highlighted –OH group—are shown, which adds to the difficulty for models trying to understand the table.
3.3.3 Reasoning Questions
Reasoning questions are designed to further analyze and infer information from the data presented in the table within the image. They require a comprehensive understanding to make informed judgments. These questions include: 1) Numerical and statistical reasoning, 2) Trend and change analysis, 3) Multi-hop logical reasoning, and 4) Domain-specific reasoning. Specific details are provided in the Appendix. Since the data in chemical tables is often tightly linked to specific experimental conditions, molecular structures, and graphical annotations, reasoning questions not only assess the ability to comprehend explicit information but also place greater demands on domain knowledge, the integration of information across rows and columns, and the ability to reason using visual symbols.
3.3.4 Data Filtering
Diversity. We identified questions with overly repetitive structures and semantically similar phrasing to encourage a broad range of question formulations. These were rewritten using GPT-4.1 with prompt templates provided in the Appendix. We selected algorithms that maximised the semantic distance from the original question, enhancing the linguistic and structural diversity of the dataset.
Difficulty. We implemented a filtering strategy to ensure the dataset poses a meaningful challenge. We first conducted a single-pass QA evaluation using the Qwen-2.5-7B model [34]. For each question, we recorded whether the model was able to produce the correct answer on the first attempt. Questions that were answered correctly in one try were deemed too simple. To filter out these low-difficulty samples, we randomly discarded some of them. This approach allowed us to enrich the dataset with more challenging examples that were better suited for evaluating advanced reasoning capabilities.
4 Experiments on Table Recognition
4.1 Experimental Setup
Evaluation Metrics.
To evaluate the performance of the table recognition task, we adopt the improved similarity metric based on tree edit distance (TEDS) [21] along with the TEDS-structure indicator. Specifically, the table content may contain molecular graphs in chemical scenarios. Interpreting these molecular graphs using the simplified molecular input line entry system(SMILES) can result in structural isomorphism, where different representations or atom orders may correspond to the same molecule, making it inappropriate for TEDS to utilise normalised edit distance as a measure of cell content recognition accuracy. Therefore, for cells containing chemical molecular graphs, we replace the normalised edit distance with the Tanimoto coefficient [35] to accurately assess the performance of table recognition. We use accuracy (ACC) as the evaluation metric for fine-grained retrieval experiments.
Baselines.
We evaluate a diverse set of MLLMs. For open-source models, we test InternVL3-78B [36], Llama-3.2-90B [37], and Qwen2.5-VL-72B [38]. We also evaluate the following proprietary models: Gemini-2.5-Flash [39], GPT-4.1, GPT-4.1-mini [30], and Claude-3-7-Sonnet [40].
Figure 3: Comparative Evaluation of MLLMs and DECIMER for Molecular Formula Recognition from Real-World and Synthetic Chemical Diagrams.
[…]
Model Category Model Name Table Recognition Value Retrieval Position Retrieval
TEDS-Struct* TEDS* ACC ACC
Proprietary Claude-3-7-Sonnet 92.58 85.40 33.89 53.06
GPT-4.1 95.48 88.93 29.60 49.49
Gemini-2.5-Flash 95.91 88.29 29.19 36.92
GPT-4.1-mini 95.25 87.50 17.01 35.16
Open-Source Qwen2.5-VL 93.12 89.45 31.72 38.35
InternVL3 94.40 86.06 29.58 33.91
Llama-3.2 93.15 87.46 29.30 32.85
4.2 Experimental Results
Main Results Analysis.
We evaluate the table recognition performance using an improved similarity metric based on TEDS, along with the TEDS-structure indicator. The results are in Table 3 and key findings are listed as follows.
(a) Small performance gap between open-source and proprietary models. Although a performance gap between open-source and proprietary models still exists, both achieve promising results in table recognition. For example, Gemini-2.5-Flash achieves 95.91 on TEDS-Struct and 88.29 on TEDS, while the open-source Qwen2.5-VL also performs competitively, scoring 93.12 on TEDS-Struct and 89.45 on TEDS. This shows that both model types possess strong capabilities in table understanding and structure reconstruction.
(b) Poor fine-grained retrieval across all MLLMs. All models show weak performance on fine-grained retrieval tasks, such as locating cell content by row-column positions or inferring positions from content. As shown in Table 3, even Claude-3.7-Sonnet, the best-performing model, only achieves an accuracy of 33.89 on value retrieval, with others performing worse. This highlights ongoing challenges in achieving precise fine-grained alignment in current MLLMs.
(c) Molecular formulas pose a key recognition bottleneck. Recognition performance declines as the number of molecular formulas in tables increases. In contrast, accuracy improves significantly when molecular formulas are absent. Specifically, models perform notably worse on chemical tables with many molecular structures than on those with plain text or simple layouts. This indicates that molecular formulas remain a key bottleneck for current models and require further optimization to improve recognition accuracy.
Figure 4: Impact of molecular formula complexity on table recognition.
Analysis of Molecular Formula Recognition.
We evaluated MLLMs on molecular formula recognition from real-world academic papers and synthetic diagrams. As shown in Figure 3, MLLMs can identify and convert molecular structures into chemical formulas, but their performance on real-world diagrams is lower than on synthetic ones, highlighting the impact of data diversity and quality on model robustness. Although MLLMs possess some chemical domain knowledge, their accuracy and reliability remain significantly inferior to specialized models (DECIMER[41]). This gap underscores the need for further improvements to enable MLLMs to effectively handle domain-specific tasks in advanced scientific applications.
Analysis of Chemical Representation Recognition.
We examine whether current MLLMs treat chemical representations (e.g., molecular formulas and structural diagrams) as noise when processing tables, as shown in Figure 4. Results indicate that even when molecular structures appear only in the surrounding context—not within the table itself—their presence still degrades performance. This suggests that chemical symbols can interfere with the model’s overall understanding. Our findings confirm that MLLMs struggle to parse and integrate multimodal chemical information, revealing a key limitation in chemical table understanding and an important direction for future improvement.
5 Experiments on Table Understanding
5.1 Experimental Setup
Evaluation Metrics.
For tasks requiring descriptive answers, we use edit distance [42] as the primary metric to assess recognition accuracy. For more open-ended question answering (QA) tasks, we adopt a binary evaluation strategy powered by GPT-4.1-nano [30], which classifies each response as either correct or incorrect [43, 44, 45],. The overall performance is then quantified by computing the accuracy (ACC) based on these binary classifications.
To ensure the reliability of this automated evaluation process, we randomly sampled 20% of the QA instances for manual verification. Human annotators reviewed the model outputs against reference answers and provided independent judgments. We then calculated the agreement rate between human evaluations and the binary assessments produced by GPT-4.1-nano . Implementation details are provided in the Appendix.
Baselines.
We evaluate a comprehensive set of general-purpose multimodal large language models (MLLMs), including both open-source and proprietary systems. The open-source models evaluated are InternVL3-78B (InternVL) [36], Llama-3.2-90B (Llama) [37], and Qwen2.5-VL-72B (Qwen-VL) [38]. Among the proprietary models, we assess Gemini 2.5 Flash (Gemini) [39], GPT-4.1, GPT-4.1 mini (GPT-mini) [30], and Claude-3.5-Sonnet (Claude) [40]. To provide an upper-bound reference, we also include a human performance baseline (Human), which consists of answers manually provided by annotators on a representative subset of the evaluation data.
Table 4: Performance comparison of MLLMs on descriptive and reasoning tasks with human in chemical table understanding. Overall, MLLMs perform impressively but are slightly outperformed by humans in complex tasks. We denote the best score in blue, and the second-best score in green.
Question Type GPT-mini GPT-4.1 Claude Gemini Qwen-VL InternVL Llama Human
     Descriptive Questions
Element Description Table Dimensions 68.31 73.07 75.39 73.10 71.42 70.50 70.27 -
Title Description 84.79 87.31 81.03 81.64 83.18 84.35 85.50 -
Annotation Description 81.86 92.94 68.93 81.12 90.23 73.27 81.39 -
Molecular Recognition 37.62 42.49 52.71 28.47 14.14 11.63 18.50 -
Visual Description Bold Description 35.88 44.27 50.38 41.22 38.93 32.82 52.73 98.99
Color Description 54.92 56.48 50.38 41.22 58.55 49.74 58.03 97.73
     Reasoning Questions
Domain- Specific QA Benzene Rings Count 49.51 52.31 62.83 75.32 59.61 47.66 21.97 94.98
Yield and Conditions 89.97 89.14 89.42 93.87 85.24 74.93 82.13 93.61
Function Based 20.78 37.30 62.06 71.70 35.37 30.23 25.83 89.27
Numerical Statistics Value Comparison 78.40 91.80 93.60 98.40 94.40 67.14 81.45 100.00
Find Min/Max 79.44 85.85 94.39 83.18 94.39 60.94 60.95 100.00
Calculate Sum 32.63 58.33 53.68 64.21 32.63 24.07 34.12 100.00
Calculate Average 22.52 44.87 66.67 72.97 26.13 24.19 33.33 98.20
Logical Reasoning Trend Analysis 75.13 81.87 83.94 88.60 74.61 76.34 55.96 98.45
Multi-hop Retrieval 83.55 84.87 88.16 87.94 82.89 80.20 81.48 98.67
5.2 Experimental Results.
We compare our method against a set of representative MLLMs on table understanding task in Table 4. Below are our main findings:
(a) General reasoning strength, numerical weakness. MLLMs achieve high accuracy (ACC > 0.8) on tasks such as trend analysis, value comparison, and finding min/max values. Specifically, Gemini scores 98.40 on value comparison and 88.60 on trend analysis. This indicates strong capability in capturing basic quantitative and logical patterns from tables. However, performance drops significantly on arithmetic-heavy tasks like calculating the sum or average, where Gemini only reaches 64.21 and 72.97, respectively, highlighting limitations in calculation-intensive reasoning.
(b) Descriptive tasks outperform domain-specific chemistry QA. MLLMs show strong performance in general descriptive tasks. GPT-4.1 achieves 92.94 on annotation description. In contrast, accuracy plummets on domain-specific chemistry questions. Gemini scores only 75.32 on benzene ring counting and 71.70 on function-based QA. This reveals limitation of current models in handling specialized chemical content, especially when integrating domain knowledge with visual input.
(c) Visual style interpretation remains challenging. Tasks that rely on recognizing formatting cues — such as bold or color highlights — yield consistently low accuracy (e.g., Llama scores 52.73 on bold description and 58.03 on color description). Even top-performing models struggle to interpret these stylistic features embedded within structured data, suggesting that visual structure understanding is still an underdeveloped capability in current MLLMs.
(d) Closed-source models dominate complex and domain tasks. Closed-source models like Gemini and GPT-4.1 consistently achieve top results across most categories — Gemini leads in value comparison (98.40), trend analysis (88.60), and several domain-specific tasks. In contrast, open-source models like InternVL and Llama lag behind, particularly in complex reasoning and chemical QA, indicating that current open models are not yet competitive in specialized, vision-language reasoning domains without further fine-tuning or domain adaptation. Additionally, despite impressive performance, these models still fall slightly short of human capabilities in complex tasks.
5.3 Analysis of Multimodal Input and Model Behavior
Impact of Input Modality.
Figure 5: Accuracy Comparison of InternVL3-78B on Chemistry Table Understanding Tasks Across Input Modalities
Model Missing Ambiguity
Col/Row Style
Llama 66.43 77.31 80.08
GPT-mini 70.29 35.64 70.36
InternVL 74.84 46.15 62.80
Qwen-VL 75.51 76.27 97.64
GPT-4.1 84.02 81.21 79.69
Gemini 92.50 73.06 93.67
Claude 97.12 90.68 98.23
Table 5: Model performance on unanswerable question categories.
In Figure 5, we evaluated three input modalities—Text QA (HTML), VQA (Image), and Hybrid QA (Hybrid)—to assess how different formats affect model performance in answering chemistry-related questions. Experimental results across tasks such as Yield and Conditions, Function Based show that Hybrid QA achieves the highest accuracy by combining textual and visual inputs, enabling a more comprehensive understanding of complex chemical structures. Text QA outperforms VQA, as converting images to text improves interpretability, although it may introduce errors due to information loss. In contrast, VQA struggles with detailed visual content, leading to higher error rates. These findings suggest that hybrid input strategies are most effective for enhancing performance, while careful handling of text conversion remains essential.
Impact of Unanswerable Questions on Model Behavior.
We examine how advanced MLLMs handle unanswerable questions by refraining from responding, as shown in Table 5. This occurs when questions exceed model capabilities or lack context, which we classify into three types: non-existent content, missing format/style, and ambiguity. Our results show that leading models can effectively determine when not to answer by using contextual understanding and reasoning, reflecting a form of self-awareness that avoids misinformation. In contrast, smaller models often fail to recognize such cases, producing incorrect or irrelevant answers. This underscores the importance of model scale and training quality for reliable and trustworthy MLLMs in question-answering and prompt engineering.
Figure 6: Case study of answering descriptive and reasoning questions with different MLLMs.
Case Study Analysis.
To evaluate the capabilities of MLLMs in handling descriptive and reasoning questions, we conducted a case study in Figure 6. The results reveal that MLLMs struggle significantly with complex reasoning tasks. Although proprietary models outperform open-source models in descriptive tasks, both exhibit notable limitations in accuracy and consistency, which highlights the challenges current MLLMs face in performing deep reasoning, underscoring the need for further advancements in multimodal reasoning capabilities.
6 Conclusion
In this work, we introduced ChemTable, a large-scale dataset and benchmark designed to evaluate multimodal large language models (MLLMs) on recognition and understanding tasks involving chemical tables. By curating over 1,300 real-world chemistry tables and annotating them with domain-specific metadata and question-answering tasks, ChemTable captured the multimodal, symbolic, and semantic challenges unique to the field. Our comprehensive evaluation revealed significant performance gaps between current MLLMs and human-level capabilities, particularly in domain-specific reasoning and molecular recognition. We believed that this dataset and benchmark would facilitate future advancements in multimodal scientific analysis and understanding.
Appendix
Appendix A Fine-Grained QA Behavior Analysis
This section presents detailed analyses of how reasoning complexity and query directionality affect the accuracy and robustness of multimodal models in scientific table question answering.
A.1 Effect of Question Hops on Answer Accuracy
Table 6: Performance of Multimodal Models on Multi-Hop QA Tasks Categorized by Hop Count.
Model Hop Count Overall
2 3 4
Claude 91.29 84.98 70.70 88.16
Gemini 90.86 86.98 69.14 87.94
GPT-4.1 87.72 79.71 71.78 84.87
GPT-mini 87.68 78.43 61.65 83.55
Qwen-VL 87.29 82.81 52.84 82.89
Llama-3.2 84.59 80.18 61.78 81.48
InternVL 83.58 78.24 59.47 80.20
To understand the impact of reasoning complexity on model performance, we conduct a fine-grained analysis of multi-hop question answering, using hop counts of 2, 3, and 4 as indicators of increasing logical depth. As shown in Table X, all models exhibit a monotonic decline in performance with increasing hop counts. For instance, Claude drops from 91.29% on 2-hop questions to 70.70% on 4-hop, while InternVL declines more sharply from 83.58% to 59.47%. This trend reflects a consistent rise in difficulty as models are required to perform more compositional and contextual reasoning over tabular data.
In the context of ChemTable, where tables encode dense, multimodal chemical knowledge—including symbolic notations, visual molecular structures, and complex reaction dependencies—multi-hop questions often require the integration of spatial, textual, and domain-specific knowledge. For example, answering a 4-hop question might involve comparing reaction conditions across multiple rows and applying chemistry knowledge such as identifying functional groups or evaluating yields under varying catalysts.
Our results show that stronger models such as Claude and Gemini maintain significantly higher accuracy on high-hop questions compared to smaller or open-source models (e.g., Qwen-VL or InternVL). This growing performance gap at higher hop levels suggests that complex multi-step reasoning tasks amplify the differences in model capabilities.
A.2 Asymmetry in Condition-Yield Table Reasoning
Figure 7: Accuracy of Multimodal Models in Answering Questions Given Conditions vs. Inferring Conditions from Outcomes.
Directionality of query plays a significant role in multimodal table question answering. In this additional experiment, we evaluate four state-of-the-art multimodal large language models (MLLMs) on two complementary QA settings: (1) Forward Prediction, where the model predicts reaction outcomes (e.g., yield) based on given conditions; and (2) Inverse Prediction, where the model must infer conditions from known outcomes. As shown in Figure 7, all evaluated models demonstrate higher accuracy on the forward task compared to the inverse one. For instance, GPT-4.1 achieves 91.74% on forward prediction but drops to 86.82% on the inverse. This performance gap suggests that MLLMs are better aligned with natural scientific reasoning—where outcomes are typically deduced from conditions—than with the reverse logic. It also reflects an asymmetry in learned representations: while models can synthesize output from structured inputs effectively, they struggle more when tasked with deducing structured inputs from outcomes, which often requires multi-hop or abductive reasoning. These findings reveal a fundamental challenge for scientific understanding in reverse reasoning settings and highlight the need for targeted training strategies to enhance backward inference in MLLMs.
Appendix B Correlation Between Response Length and Answer Correctness
Figure 8: Accuracy and Sample Distribution by Thinking Length Across Models.
To investigate how reasoning depth affects model performance, we analyzed the relationship between response (or "thinking") length and accuracy on Function-Based QA tasks. We evaluated three representative Multimodal Large Language Models—GPT-4.1, Gemini-2.5-Flash, and InternVL—by binning their outputs based on token length and computing corresponding accuracies and sample counts (Figure 8). This setting helps reveal whether longer responses lead to more accurate reasoning. As shown in the results, accuracy does not monotonically improve with longer thinking length. Instead, models tend to achieve peak performance at moderate length ranges. Beyond these ranges, accuracy either plateaus or decreases, likely due to verbosity, sample counts drop sharply at extreme lengths, limiting statistical confidence in those bins. Overall, the results suggest that effective reasoning often corresponds to an optimal response length—neither too short nor excessively long.
Appendix C Effect of Chain-of-Thought Reasoning
Figure 9: Impact of Chain-of-Thought (CoT) Reasoning on Question Answering Accuracy Across Different Task Types.
We conducted an ablation study using GPT-4.1 to evaluate the impact of Chain-of-Thought (CoT) prompting on multimodal question answering over chemical tables. Specifically, we compared model performance with and without CoT reasoning across four representative question types. As illustrated in Figure 9, the removal of CoT resulted in a consistent drop in accuracy, particularly for reasoning-oriented tasks. While annotation description—largely descriptive in nature—showed minimal change (92.94% with CoT vs. 92.35% without), substantial declines were observed in function-based questions (37.30% → 24.64%), summation (58.33% → 42.18%), and trend analysis (81.87% → 76.53%). These results highlight the critical role of CoT prompting in enhancing the model’s ability to perform step-by-step reasoning and complex inference. Overall, the findings underscore that even high-capacity models like GPT-4.1 benefit significantly from structured reasoning guidance, particularly in reasoning QA scenarios.
Appendix D Specifications of Annotation Format and Procedure for Table Recognition
Figure 10: Structured Annotations of a Chemical Table into Title, Reactions, Substances, Table, and Annotations Components.
D.1 Annotation Format
During the table image annotation phase, we conducted detailed annotations of chemical reaction tables to facilitate the structured extraction and downstream machine learning tasks. Specifically, we divided each chemical table into five distinct components: Title, Reactions, Substances, Table, and Annotations. The structure of each component is systematically defined to ensure consistency and interpretability across the dataset, as illustrated in Figure 10.
On the left side of Figure 10, we present a typical chemical reaction table that has been comprehensively annotated. This includes:
  • Title: The title of the table, annotated with a bounding box, unique index, and its content. Indicate the subject or purpose of the table, such as the optimization of reaction conditions.
  • Reactions: Schematic representations of chemical reactions including clearly separated reactants, products, and conditions. Each reaction entry is annotated with a unique identifier, text content, and a bounding box. Additionally, the involved chemical entities are linked to corresponding substance entries.
  • Substances: All unique chemical structures in the table are annotated with bounding boxes and identification text. Each structure is mapped to its SMILES representation and identifier, ensuring its traceability and enabling computational applications.
  • Table: The core tabular data, where each cell is annotated not only with a bounding box and textual content but also with logical coordinates indicating its specific row and column position in the table. Furthermore, chemical substances mentioned within the table cells are linked to the corresponding SMILES representations via a mapping mechanism.
  • Annotations: Footnotes or explanatory text found below or around the table are included as annotations. These provide essential context such as experimental conditions or special notes on reagents and are annotated with bounding boxes and textual content.
On the right side of Figure 10, the data structure for each annotation component is illustrated using JSON-style schema definitions. This schematic defines the internal representation of each component in the dataset, including keys like "bbox", "text", "maps", "start_row", and "start_col", enabling precise spatial and logical referencing within the table.
This annotation scheme ensures that both the visual layout and the semantic structure of chemical tables are faithfully captured, which is crucial for downstream applications such as automated chemical information extraction and chemical literature understanding.
D.2 Fine-Grained Text Annotation Rules
To further enhance the interpretability and structured utility of the dataset, we adopted a standardized markup protocol for annotating textual elements within the table components:
[…]
  • Subscripts and Superscripts: Chemical text annotations follow the convention of using the caret symbol (^) for superscripts and the underscore (_) for subscripts.
  • Text Formatting: For stylistic features embedded in the body of the table (excluding titles and substance labels), the following LaTeX-style conventions are used:
[…]
These conventions are designed to faithfully capture the nuanced visual and semantic cues present in scientific tables, which are critical for tasks involving automatic parsing, entity recognition, and domain-specific layout understanding.
D.3 Annotation Procedure
The annotation process was carried out in three well-defined phases to ensure both coverage and precision across diverse types of chemical reaction tables.
Phase I: Data Collection and Categorization.
We began by collecting and curating table images from peer-reviewed chemical literature published over the past decade. Specifically, we sourced documents from high-impact journals such as ACS Catalysis, Journal of the American Chemical Society, Chem, Angewandte Chemie International Edition, and Angewandte Chemie. From these publications, we systematically extracted regions explicitly labeled as tables based on captions, figure titles, or context cues. Following extraction, each table image was manually categorized into one of six functional types based on its primary purpose: (1) Reaction Condition Optimization Tables, (2) Substrate Scope Tables, (3) Chemical Structure Information Tables, (4) Reaction Feature Data Tables, (5) Property/Outcome Comparison Tables, and (6) Statistical Summary Tables. Among these, optimization and substrate scope tables comprised the majority (over 50%), while the remaining types each accounted for more than 10% of the dataset.
Phase II: Coarse-Grained Annotation.
In the second phase, we performed coarse-grained annotations to establish the structural foundation of each table image. This included identifying and labeling five primary components — Title, Reactions, Substances, Table, and Annotations. For each component, bounding boxes and textual transcriptions were annotated. For instance, the Title region was delineated and transcribed to capture the overarching context of the table. Reaction schemes were segmented and annotated as Reactions, while molecular structures embedded in the table were marked as Substances. The main data matrix was annotated under the Table component, and any surrounding descriptive notes or footnotes were included under Annotations. This phase focused on accurately demarcating high-level semantic units to support later fine-grained processing.
Phase III: Fine-Grained Annotation.
In the final phase, we applied fine-grained annotations to the core tabular content and reaction schematics. For the Table component, we annotated each individual cell with its bounding box, textual content, and logical position (row and column indices). If a cell referenced chemical entities, we linked the corresponding text or image to its associated SMILES identifier using a predefined mapping. Similarly, in the Reactions component, we annotated and linked specific chemical species — such as reactants, products, catalysts, or solvents — and reaction conditions (e.g., temperature, time, yield) to structured representations, enabling both visual and semantic disambiguation.
Following this three-stage annotation pipeline, we constructed a high-quality dataset comprising 1,500 fully annotated chemical table images. This dataset preserves both the visual layout and the underlying chemical semantics, laying a robust foundation for downstream tasks including machine learning model training, automated reaction information extraction, and chemical table understanding.
Appendix E Algorithm for Converting Annotations to HTML
Since we only give the logical coordinate annotations of the tables, they cannot be directly used to calculate the TEDS metrics. To address this, we use the following pseudocode to convert the logical structure into markup sequence format. Firstly, we divide the entire conversion process into two stages. As shown in Algorithm 1, the first stage involves a preliminary preprocessing of the logical location information, associating the logical positions with the corresponding cell content and storing them in a tabular data matrix.
Algorithm 1 From Logical Location to Tabular Data Matrix
Input: cells = { $C_{1},C_{2}\dots,C_{K}$
Output: table
1:max_row $\leftarrow$ maximum value of ’end_row’ in cells
2:max_col $\leftarrow$ maximum value of ’end_col’ in cells
3:Initialize table as a array with dimensions (max_row, max_col)
4:for cell in cells do
5:    start_row, end_row, start_col, end_col, content $\leftarrow$ cell
6:    rowspan $\leftarrow$ 1 $+$ end_row -- start_row
7:    colspan $\leftarrow$ 1 $+$ end_col -- start_col
8:    for row = start_row to end_row do
9:         for col $\leftarrow$ start_col to end_col do
10:             if row ==absent$=$ = start_row and col ==absent$=$ = start_col then
[…]
12:                 "colspan": colspan, "content": content }
13:             else
14:                 table[row][col] $\leftarrow$ "merged"
15:             end if
16:         end for
17:    end for
18:end for
In the second stage, we traverse the tabular data matrix row by row, gradually converting the stored logical information and cell content into a mark-up sequence, eventually generating the conversion result. The specific implementation is detailed in Algorithm 2. The corresponding source code is available in our GitHub repository.
Algorithm 2 From Tabular Data Matrix to Markup Sequence
Input: table
Output: markup
1:Initialize markup $\leftarrow$ "<table>"
2:for row in table do
3:    markup += "<tr>"
4:    for cell in row do
5:         if cell == "mergedthen
6:             continue
7:         else
8:             rsp = cell["rowspan"]
9:             csp = cell["colspan"]
10:             markup += <td rowspan $=rsp\text{ colspan}=csp\texttt{>}$
11:             markup += cell["content"] + "</td>"
12:         end if
13:    end for
14:    markup += "</tr>"
15:end for
16:markup += "</table>"
Appendix F Workflow for Question Annotation
The process of question annotation in ChemTable is structured into four complementary stages: rule-based automatic generation, LLM-assisted synthesis, manual refinement, and domain-specific function-based annotation. Together, these stages ensure that the resulting question-answer pairs are both scalable and chemically meaningful. Each stage contributes a distinct layer of complexity and reasoning depth—from surface-level descriptions to functionally grounded scientific inquiry—allowing ChemTable to comprehensively cover the diverse landscape of tabular chemical data.
F.1 Rule-Based Automatic Annotation
The first stage of question annotation is grounded in the structured annotations obtained from the preceding Table Recognition phase. Specifically, we utilize layout and semantic information such as table title, annotation blocks, cell types, and molecular elements. A set of deterministic scripts are designed to automatically generate descriptive question-answer (QA) pairs based on predefined heuristics. These rules capture basic factual and metadata-related queries, such as:
  • Table dimensions (e.g., number of rows and columns),
  • Title description (extracting the title),
  • Annotation interpretation (e.g., footnotes or notes),
  • Molecular recognition (e.g., identifying the molecular structure type in a cell).
This rule-based method ensures high coverage and consistency across common question types, especially those targeting descriptive understanding without the need for deep inference.
F.2 LLM-Assisted Generation for Simple Reasoning
For numerical and statistically oriented reasoning tasks, we adopt a semi-automatic pipeline leveraging large language models (LLMs), specifically GPT-4.1. The question generation follows a prompt-based paradigm where we input the HTML representation of the table, the table image, and a specified reasoning type (e.g., comparison, summation) into a structured prompt template (see Section L for details). The model then outputs a set of QA pairs. The LLM is guided to focus on quantifiable patterns and basic statistical operations, such as:
  • Value Comparison (e.g., comparing yields across rows),
  • Find Min/Max (e.g., identifying the entry with the highest selectivity),
  • Calculate Sum (e.g., summing up yields in a column),
  • Calculate Average (e.g., computing the mean conversion).
F.3 Manual Annotation for Complex Reasoning
For questions involving complex domain-specific logic or requiring visual-semantic integration—such as multi-hop reasoning, ambiguous references, or molecular structure interpretation—we rely on manual annotation. Annotators use an internal tool, LabelStudio (see Section H), where they are presented with the image of the table, metadata, and a set of predefined question types.
Human annotators are instructed to create diverse and challenging questions that demand:
  • Domain knowledge (e.g., understanding catalyst-function relationships),
  • Logical inference (e.g., combining footnotes with table entries),
  • Visual decoding (e.g., counting specific molecular motifs like benzene rings).
To ensure quality, each question undergoes two rounds of review: first by MLLMs validation and then through another annotator validation. Annotators are also encouraged to include unanswerable questions–caused by non-existent content, missing format/style, and ambiguity—to reflect real-world data imperfections to further test model robustness.
F.4 Domain-Specific Question Annotation: Function-Based QA
While prior work in scientific table QA has largely focused on fixed question templates—such as yield estimation, comparison, or description—chemical tables exhibit significantly broader functional diversity. In ChemTable, we identify a substantial subset of tables whose structure and purpose deviate from conventional paradigms. These include, but are not limited to, substrate screening matrices, catalyst performance evaluations, structure-property tables, and experimental condition explorations. Unlike standard output-driven tables (e.g., yield-focused), these tables encode specific scientific functions that are often implicit and domain-specific.
To address this, we introduce a novel annotation paradigm termed Function-Based QA, which aims to capture questions grounded in the functional roles of tables within experimental workflows. This process is semi-automated and comprises the following pipeline:
  1. 1.
    Function Summary Generation. We begin by prompting GPT-4.1 with the full HTML representation and image of a table, guiding it to generate a concise natural language summary that articulates the table’s experimental function, purpose, or analytical focus.
  2. 2.
    Function-Aligned Question Generation. Based on the generated summary and table contents, we prompt GPT-4.1 and Claude to produce candidate QA pairs that probe aspects closely tied to the table’s described function. These include nuanced inquiries such as the effect of a specific ligand under a fixed condition, rationale for substrate ordering, or interpretation of experimental design variables.
  3. 3.
    Validation via Multi-Round QA. To verify correctness and answerability, each candidate question undergoes three rounds of independent answering by GPT-4.1 and Claude-3.7-Sonnet. If all answers are consistent and correct across rounds, the question is accepted as valid. If discrepancies arise, human annotators inspect the question-answer pair for correctness and revise or discard as needed.
This annotation strategy enables ChemTable to extend beyond rigid QA formats, capturing richer scientific inquiry styles that reflect how chemists interpret and utilize tabular data. The resulting Function-Based QA subset significantly improves the coverage of the benchmark on real-world analytical reasoning.
Appendix G Consistency Analysis: Verification of Human vs. MLLM
To validate the reliability of our automatic QA evaluation pipeline powered by GPT-4.1-nano, we randomly sampled 20% of the QA instances for manual verification. Two human annotators with chemistry backgrounds independently judged whether model-generated answers were correct, referencing the original table content and gold answers. We calculated agreement accuracy between human and automated judgments using simple percentage overlap:
Agreement $\text{Agreement Rate}=\frac{|\mathcal{J}_{\text{human}}\cap\mathcal{J}_{\text{GPT}}|}{|\mathcal{J}_{\text{sampled}}|},$ (1)
where $\mathcal{J}_{\text{human}}$ and $\mathcal{J}_{\text{GPT}}$ denote the sets of instances labeled as correct by human annotators and GPT-4.1-nano, respectively, and $\mathcal{J}_{\text{sampled}}$ is the total set of sampled QA instances used for verification. The comparison between human annotations and GPT-4.1-nano’s binary classifications showed a high agreement rate of 96.8% overall. Agreement was particularly strong for descriptive and numerical tasks. These results confirm that GPT-4.1-nano provides a reliable and scalable approximation of human judgment for most evaluation scenarios in ChemTable, although human review remains valuable for borderline or high-complexity cases.
Appendix H Screenshots of the Annotation Interface
Figure 11: LabelStudio Dashboard View Showing Task Categories for Question-Answer Annotation.
We utilize LabelStudio as the primary platform for all data annotation tasks. The system is deployed on our internal computing clusters, allowing annotators to securely access the annotation interface via SSH forwarding. As shown in Figure 11, annotators begin by selecting a designated task from the project dashboard, where each task corresponds to a specific QA category (e.g., Yield and Conditions, Visual Reasoning, Benzene Rings).
Once a task is selected, the annotator is presented with the annotation interface (Figure 12), which displays the target table image at the top. Annotators must first select a suitable question type label. Based on the selected label, they are required to design a corresponding question-answer pair by analyzing the table content. To aid consistency, a reference panel on the right side of the interface provides example questions tailored to the selected type.
Figure 12: Annotation Interface for Generating QA Pairs from Table Content: Annotators Select a Question Type, Input the Question and Answer, and Refer to Provided Examples on the Right.
We encourage free-form question formulation to improve diversity. However, we enforce a strict constraint: all answers must be either directly found in the table or logically inferable from it without requiring specialized chemical knowledge. This ensures that questions remain grounded and accessible. In addition, annotators are encouraged to include unanswerable questions when the table content is ambiguous or insufficient. Such cases must be clearly noted in the Remark field to support subsequent filtering or diagnostic analysis.
Appendix I Implementation Details
To ensure consistency and stability across different models during evaluation, we adopted a unified decoding configuration for all multimodal large language models. Specifically, the temperature was set to 0.0 and the nucleus sampling parameter (top-p) to 0.2, wherever supported. This configuration minimizes randomness and promotes deterministic outputs, which is essential for fair and reproducible evaluation across both table recognition and question answering tasks. This setup also facilitates meaningful comparisons between models under controlled conditions.
Appendix J Dataset Visualization by Image Type
Figure 13: Representative Examples of Six Chemical Table Types in the ChemTable Dataset.
Appendix K Taxonomy of Question Types with Representative Cases
Table 7: Representative Question Across Different Task Types for Chemical Table Understanding.
Question Type Question Case 1 Question Case 2 Question Case 3
Table
Dimensions
What is the size of the table in the picture? - -
Title
Description
What is the title of this table? - -
Annotation
Description
What are the annotations of this table? - -
Visual
Description
What is the reaction time in the row highlighted in light blue? In the catalyst column, what is the content in bold? For the row highlighted in light blue, which has a higher yield, 3a or 4a?
Benzene Rings
Count
How many molecular diagrams of benzene rings are there in the table? How many benzene rings are in the diagram? What is the proportion of substances containing benzene rings among all the substances in the table?
Yield and
Conditions
What is the reaction time when the yield is at its highest? Under the condition of 50°C, at what reaction time is the yield highest? At a temperature of 70°C and a reaction time of 30 minutes, which has a higher yield, 3a or 4a?
Function
Based
What is the yield (%) of 3f at the reaction time where the yield first drops below half of its maximum value? At which entry does the yield of 3f become less than the yield observed at 30 min, and what is the yield at that entry? Which structure has the highest number of solvent atoms, and what is that number?
Numerical
Statistics
What is the mean (average) value of I0/I? At 223 K, which is higher: the calculated $v_{a}-v_{b}$ or the observed $v_{a}-v_{b}$? What is the sum of the yields of product 2 across all solvents?
Trend
Analysis
How does the yield of 3p change with increasing reaction time? What is the trend in the third column? As k increases, what is the trend in the change of obs.$v_{a}-v_{b}$?
Multi-hop
Retrieval
Which entries in the table have a yield of 99? What is the maximum value in the U(eq) column of the table? In the figure, what is the reaction time corresponding to a 10% yield of 6c?
Appendix L Prompt Templates
L.1 Table Recognition Prompt Templates
Prompt 1 is designed to evaluate a model’s ability to extract and reconstruct the structural layout of a table from an input image. The instruction explicitly requests the HTML representation of the table using only five basic tags: <table>, <thead>, <tbody>, <tr>, and <td>. To ensure a semantically accurate output, the prompt emphasizes the separation of the table header and body using <thead> and <tbody> tags, respectively. This task tests the model’s understanding of both visual layout and hierarchical table semantics without reliance on style or advanced formatting. It serves as a foundational prompt for assessing table structure recognition capabilities.
Figure 14: HTML Table Structure Identification from Images.
Prompt 2 focuses on cell-level retrieval by requiring the model to locate the exact position of a cell within a table image based on a given content string. The model must identify the table structure, search for the specified cell content, and return the corresponding row and column indices in a structured JSON format.
Figure 15: Locate Cell Position by Content in Table Images.
Prompt 3 evaluates the model’s ability to locate and extract the textual content of a specific cell based on given row and column indices within a table image. The coordinates are one-indexed and cover both the header and body of the table. The model must parse the structure visually, identify the correct cell, and return its string content in a JSON object.
Figure 16: Retrieve Cell Content by Position in Table Images.
Prompt 4 targets the task of molecular recognition by instructing the model to identify molecular structures in a given image and convert them into SMILES (Simplified Molecular Input Line Entry System) format. This task evaluates a model’s capacity for visual parsing of chemical diagrams, structural interpretation, and chemical knowledge alignment. The expected output is a valid SMILES string encapsulated within <smiles> tags, ensuring format consistency.
Figure 17: Molecular Recognition and SMILES Conversion from Images.
L.2 Table Recognition Prompt Templates
Prompt 5 evaluates a model’s ability to extract the title text from a scientific document image. The expected output is a JSON object containing both the extracted title and a brief explanation outlining the reasoning process.
Figure 18: Extract Document Title from Image.
Prompt 6 asks the model to answer a question by reasoning over an HTML-rendered table. Given the structured table content and a specific question, the model must provide the answer to the question.
Figure 19: Answer Table-Based Question from HTML Structure.
Prompt 7 requires the model to identify the total number of rows and columns in a table image. The output is formatted as a JSON object with a reasoning chain and a count of rows and columns. The task evaluates the model’s capability in structural table parsing and its consistency in counting visual elements in scientific layouts.
Figure 20: Retrieve Table Dimensions from Image.
Prompt 8 challenges the model to answer a natural language question based on a table image. This tests the model’s vision reasoning ability by requiring joint understanding of the image content and question intent.
Figure 21: Visual Table Question Answering.
Prompt 9 evaluates the model’s ability to answer questions using both a table image and its text. By combining unstructured (visual) and structured (HTML) inputs, this prompt tests how effectively the model integrates both modalities to improve accuracy and handle noise or ambiguity in either format.
Figure 22: Multimodal Table QA with HTML and Image Inputs.
Prompt 10 asks the model to produce five question–answer pairs in statistical categories (e.g., max, sum, mean, compare) based on a table image and its HTML content.
Figure 23: Generate Statistical Questions and Answers from Table.
Prompt 11 requires the model to assess whether a given answer correctly responds to a question. It must return a binary decision (“correct” or “incorrect”) and explain its reasoning.
Figure 24: Answer Evaluation and Judgement Prompt.
Prompt 12 tasks the model with summarizing the content and purpose of a scientific table, based on its HTML structure and visual appearance. The expected summary should concisely capture the data’s meaning, key variables, and scientific implications.
Figure 25: Table Summarization for Scientific Contexts.
Appendix M Discussion of Limitations
This work focuses on evaluating general-purpose multimodal large language models on chemical tables but does not involve domain-specific pretraining or instruction tuning. As a result, current models may lack sufficient inductive bias for accurately interpreting chemical symbols, terminology, and molecular graphics. Future research could explore domain adaptation strategies such as finetuning (e.g., LoRA) to enhance models’ domain sensitivity.
Moreover, our evaluation isolates table images from surrounding textual context (e.g., experimental descriptions, chapter text), which often serve as essential semantic cues in scientific reading. Incorporating such context may improve performance on complex reasoning. Finally, the current benchmark emphasizes single-table understanding, whereas scientific synthesis often requires reasoning across multiple tables or documents. Extending the benchmark to multi-table or multi-document settings remains an important direction for future work.
Appendix N Potential Impacts
The ChemTable benchmark has the potential to positively influence both the scientific and machine learning communities by facilitating more accurate and structured extraction of domain-specific knowledge from chemical literature. This capability could accelerate AI-assisted scientific discovery, streamline the interpretation of experimental data, and improve accessibility for researchers engaging with complex multimodal content.
Nevertheless, there are possible negative societal impacts. Multimodal models applied to scientific documents may generate incorrect or misleading outputs, especially in high-stakes domains such as chemistry, where hallucinations or misinterpretations of molecular structures or reaction conditions could lead to unsafe conclusions. Moreover, if training data lacks diversity across chemical subfields or journal types, resulting models may exhibit biases or limited generalizability.
To mitigate these risks, we released the dataset under a permissive license and ensured that all data were sourced from publicly available and ethically compliant literature. We encourage future users to implement responsible usage policies, apply model interpretability tools, and evaluate models in safety-critical domains with domain-expert supervision.
Appendix O Copyright and Licensing
The dataset presented in this work was constructed by collecting and processing table images extracted from published scientific articles. We ensured that all source articles fall under licenses that permit such use, including CC0, CC-BY 4.0, CC-BY-SA 4.0, and CC-BY-NC 4.0. For each extracted table image, we have provided explicit attribution to the original publication, including its DOI.
The journals from which images were sourced include:
After annotation and compilation, the resulting dataset is released under the Creative Commons Attribution-ShareAlike 4.0 International (CC BY-SA 4.0) license, which permits reuse, redistribution, and adaptation, provided appropriate credit is given and any derivative works are licensed under the same terms. Our code is licensed under the Apache License 2.0. All table images are subject to the copyright terms of their original publications and publishers.